OpenStudy (cp123):

What is the second root of 27 plus the second root of 3 equal to?

OpenStudy (cp123):

I need help simplifying this one. I know 27 needs to be simplified to 3 x 3 x 3. but i dont know how to simplify it further. wouldnt it be \[\sqrt[3]{3} + \sqrt{3}\]

OpenStudy (cp123):

but im not sure how to simply further. i feel like it is \[\sqrt[4]{3}\] but im not sure :(

OpenStudy (sshayer):

\[\sqrt{27}-\sqrt{3}=\sqrt{9 \times 3}-\sqrt{3}=3 \sqrt{3}-\sqrt{3}=2\sqrt{3}\]

OpenStudy (cp123):

thanks but it is addition not subtraction

OpenStudy (sshayer):

sorry ,i mixed with another question. \[3\sqrt{3}+\sqrt{3}=4\sqrt{3}\]

OpenStudy (cp123):

thanks okay. thank you. thats what I thought but I am out of practice so i couldnt remember.